### v1

# LongViTU: Instruction Tuning for Long-Form Video Understanding

Abstract
This paper introduce LongViTU, a large-scale (~121k QA pairs, ~900h videos), automatically generated dataset for long-form video understanding. We developed a systematic approach that organizes videos into a hierarchical tree structure and incorporates self-revision mechanisms to ensure high-quality QA pairs. Each QA pair in LongViTU features: 1) long-term context (average certificate length of 4.6 minutes); 2) rich knowledge and condensed reasoning (commonsense, causality, planning, etc.); and 3) explicit timestamp labels for relevant events. LongViTU also serves as a benchmark for instruction following in long-form and streaming video understanding. We evaluate the open-source state-of-the-art long video understanding model, LongVU, and the commercial model, Gemini-1.5-Pro, on our benchmark. They achieve GPT-4 scores of 49.9 and 52.3, respectively, underscoring the substantial challenge posed by our benchmark. Further supervised fine-tuning (SFT) on LongVU led to performance improvements of $12.0\%$ on our benchmark, $2.2\%$ on the in-distribution (ID) benchmark EgoSchema, $1.0\%$, $2.2\%$ and $1.2\%$ on the out-of-distribution (OOD) benchmarks VideoMME (Long), WorldQA and OpenEQA, respectively. These outcomes demonstrate LongViTU’s high data quality and robust OOD generalizability.
Figure 1: Illustration of LongViTU.  The top row shows an example video sequence, with red boxes highlighting key clues for the posed question and yellow boxes marking objects in key frames related to the answer. The middle row emphasizes the primary advantages of our proposed LongViTU over previous datasets, along with a Q&A sample; refer to section 1 for further details. The bottom row displays predictions from canonical open-source and commercial VLMs, evaluated by GPT-4 against ground truth using our novel predefined criteria.
1 Introduction
We present LongViTU, a novel dataset for large-scale, long-form video understanding (see Figure 1). Compared to existing Video Question-Answering (VQA) datasets, LongViTU provides greater naturalness and diversity. The primary advantages over prior works are summarized below, see subsubsection 2.1.1 for a clearer view.
  • $\bullet$
    Diverse Real World Scenarios.  Existing datasets suffer from limitations in domain coverage or scene diversity, like Env-QA [17] and OpenEQA [35] rely on videos from virtual environments, which inherently introduce a substantial domain gap. Others, including EgoVQA [14], EgoTaskQA [22], EgoSchema [36], WorldQA [67], MoVQA [64], MovieChat [45], and CinePile [41], etc., despite utilizing real world videos, primarily feature limited or homogeneous scenes. In contrast, LongViTU leverages the comprehensive Ego4D dataset [19], enabling VQA tasks to encompass a broader spectrum of real-world scenarios with enhanced diversity and realism.
  • $\bullet$
    Explicit Timestamp Labels.  Previous datasets, such as Otter [25], Video-ChatGPT [34], InternVideo [51], VideoChat [28], MVBench [29], LLaVA-Video [66], and MM-Ego [60], lack explicit timestamp labels for QA-related events. This omission means the precise start and end times for each QA are undefined, despite multiple QAs being present in a single video. To address this, our hierarchical pipeline organizes video content into a tree structure, enabling QA generation at various granularities with explicit timestamps for each event. LongViTU provides precise temporal annotations for all QA events, facilitating accurate identification of key moments in ultra-long videos. This approach shifts away from turn-based QA paradigms by LLMs & VLMs, enabling streaming QA for extended video content.
  • $\bullet$
    Long Certificate Length.  In short datasets, such as NextQA [54] and ActivityNet-QA [62], the average certificate length (introduced in EgoSchema [36], which we adhere to) is typically less than 10 seconds. While longer datasets like WorldQA [67] reach average certificate length under 60 seconds, and EgoSchema [36] stays below 100 seconds. LongViTU achieving an average certificate length of 276.8 seconds (~4.6 minutes). For further statistical details, please refer to Figure 3.
  • $\bullet$
    Fine-grained Categorization.  Most datasets focus primarily on basic question types related to spatial elements (such as objects, attributes, locations, and states, etc.), lacking a detailed and structured categorization. While recent datasets like EgoTaskQA [22] and OpenEQA [35] have introduced categorized questions, their classifications remain broad. In contrast, LongViTU emphasizes the spatiotemporal dimensions, offering a fine-grained categorization that captures the intricate spatiotemporal details within video contexts. This is organized into three primary categories, along with more granular subcategories, illustrated in Figure 3(b).
  • $\bullet$
    Open-ended Precise QA.  Unlike existing datasets with multiple-choice formats or irrelevant answers, LongViTU delivers open-ended precise QAs. By integrating a self-revision mechanism, LongViTU refines QAs to ensure alignment with video content and removing redundancies.
By preserving these primary advantages during dataset construction, the final dataset comprises ~121k high-quality QA pairs within ~900 hours of videos. To the best of our knowledge, LongViTU is the first publicly available long-form video question-answeringing dataset featuring explicit QA-related timestamp annotations. Our main contributions are summarized as follows:
  • $\bullet$
    We propose a novel automatic pipeline to generate high-quality video question-answeringing data, mitigating several limitations of existing datasets: diverse real world scenarios, explicit timestamp labels, long certificate length, fine-grained categorization, and open-ended precise QA.
  • $\bullet$
    With our pipeline, we curate LongViTU, a large-scale high-quality pre-training dataset and benchmark aimed at advancing instruction tuning for long-form and streaming video understanding.
  • $\bullet$
    We conducted extensive experiments demonstrating the benefits of LongViTU for canonical open-source and commercial VLMs and providing insightful analysis.
2 The LongViTU Dataset
We developed a hierarchical pipeline for processing long-form videos by structuring the content into a tree format. This approach effectively generates QA pairs with explicit timestamps and extended certificate lengths, while preserving fine-grained spatial details and temporal relevance across diverse scales.
2.1 Dataset Pipeline
2.1.1 Stage I: Hierarchical Video Tree Construction
Frame Level.  Commencing at the frame level, we leverage InternLM-XComposer2 [13] to perform multi-frame dense captioning (sampled at 1 fps) across annotated events in the Ego4D [19]. Descriptions are structured as $\langle d_{f},t_{s}^{f},t_{e}^{f}\rangle$, containing frame-level text description $d_{f}$ and accurate timestamps $t_{s}^{f}$ and $t_{e}^{f}$ derived from Ego4D’s temporal annotations for each event provided by human annotators.
$\mathcal{F}^{k}=\langle d_{f}^{k},t_{s}^{f^{k}},t_{e}^{f^{k}}\rangle,\quad k=1,\dots,F$ (1)
where $F$ denotes the total number of sampled frames.
Event Level.  To eliminate redundant frame level text, we apply GPT-4 [1] to refine both manually annotated events from Ego4D and automated dense captions, yielding concise event level descriptions $\langle d_{e},t_{s}^{e},t_{e}^{e}\rangle$ with accurate temporal boundaries $t_{s}^{e}$ and $t_{e}^{e}$.
$\mathcal{E}^{j}=\langle d_{e}^{j},t_{s}^{e^{j}},t_{e}^{e^{j}},\{\mathcal{F}^{k}\}_{k=1}^{F}\rangle,\quad j=1,\dots,E$ (2)
where $E$ denotes the total number of events.
Table 1: Comparison with previous datasets.  The video sources for each dataset are listed under "Base", where "Collection" indicates that videos are derived from various resources. Furthermore, ✗ denotes multiple-choice answers, while ✓∗∗ indicates open-ended answers, LongViTU is the first large-scale dataset designed for long-form and streaming video understanding with explicit timestamp labels. The video durations and the number of QA pairs are approximate.
Dataset Base Scenario Open-ended Answer Fine-grained Categorization Explicit Timestamp Video Duration QAs
EgoVQA [14] IU Multiview [55] real world    ✗ 10h 600
Env-QA [17] AI2-THOR [24] virtual env       ✓∗∗ 130h 85.1K
EgoTaskQA [22] LEMMA [21] real world 15h 40K
MoVQA [64] Collection movie 50h 22K
MovieChat [45] Collection movie 160h 13K
CinePile [41] Collection movie 420h 303K
OpenEQA [35] ScanNet [11] HM3D [40] virtual env 3h 1.6K
WorldQA [67] PVSG [59] real world 10h 1K
EgoSchema [36] Ego4D [19] real world 250h 5K
LLaVA-Video [66] Collection real world 2000h 960K
MM-Ego [60] Ego4D [19] real world 3000h 7M
LongViTU (ours) Ego4D [19] real world 900h 121K
Segment Level.  GPT-4 further organizes these events into segments in the hierarchical video tree $\mathcal{T}_{\text{video}}$, merging related events into segments summarized at the segment level, represented as $\langle d_{s},t_{s}^{s},t_{e}^{s}\rangle$.
$\mathcal{S}^{i}=\langle d_{s}^{i},t_{s}^{s^{i}},t_{e}^{s^{i}},\{\mathcal{E}^{j}\}_{j=1}^{E}\rangle,\quad i=1,\dots,S$ (3)
where $S$ denotes the total number of segments.
Video Tree Formulation.  Drawing on the pipeline above, we formalize the hierarchical tree structure for long-form video content as follows:
$\mathcal{T}_{\text{video}}=\langle R,\{\mathcal{S}^{i}\}_{i=1}^{S}\rangle$ (4)
where $\mathcal{T}_{\text{video}}$ represents the hierarchical tree, with $R$ as the root node and nodes down to frames. Figure 2 illustrates the whole architecture.
2.1.2 Stage II: Long-Form QA Generation
Sliding Window.  A sliding window operation on video subtrees generates QAs by balancing spatial details and temporal sequences. A five-segment window captures long-term temporal relevance and short-term spatial features, questions are derived from the first three segments, and answers from the subsequent two. The hyperparameters are set based on the duration distribution of the Ego4D dataset and can be adjusted for specific task requirements.
QA Generation.  We generate QA pairs from subtrees within the sliding window, formalized as:
$\mathcal{QA}_{\text{video}}=\big{\langle}Q,A,\mathcal{D}_{s}\big{\rangle}$ (5)
Here, $\mathcal{QA}_{\text{video}}$ denotes the set of generated QA pairs for the video subtree $\mathcal{T}_{\text{video}}$. The notation $\mathcal{D}_{s}$ represents the selected segments via the sliding window, where each segment $\mathcal{D}_{s}=\big{\langle}d_{s}^{i},\mathcal{D}_{e}\big{\rangle}$ $(s=1,\dots,S)$ consists of event-level descriptions $\mathcal{D}_{e}=\big{\langle}d_{e}^{j},\mathcal{D}_{f}\big{\rangle}$ $(e=1,\dots,E)$. Each event further includes frame-level descriptions $\mathcal{D}_{f}=\big{\langle}d_{f}^{k}\big{\rangle}$ $(f=1,\dots,F)$.
Self-Revision.  In this stage, GPT-4 rigorously reviews the generated question-answer pairs to verify their consistency with the video content. It identifies and corrects any inconsistencies or inaccuracies, eliminates irrelevant or unnecessary information, and emphasizes the core elements of the QA pairs to mitigate redundancy and overly simplistic responses. Additional details on the prompts and human evaluation procedures for this process are provided in Appendix C.
2.2 Characteristics and Statistics
Key Advantages.  Our two-stage methodology for constructing LongViTU aligns with the hierarchical structure of the video tree, offering several advantages that significantly enhance the data quality and utility:
  • $\bullet$
    Explicit Timestamp Labels: Each tree node is annotated with precise event timestamps, enhancing the reliability of temporal analysis.
  • $\bullet$
    Long Certificate Length: Comprehensive representation of event sequences ensures robust QA generation by covering longer temporal spans.
  • $\bullet$
    Fine-Grained Categorization: By leveraging specific subtrees, the methodology optimizes input length for large language models (LLMs), enabling effective processing of extended temporal events, preservation of spatial details, and support for QA generation across detailed categories.
Figure 2: Pipeline of LongViTU.  We adopt a hierarchical pipeline that organizes video content into a tree structure, with subtrees encapsulating information at different temporal scales. This framework facilitates the generation of QA pairs with explicit timestamps, ensuring adaptability to varying contextual lengths. By summarizing content across multiple temporal levels (frame level, event level, segment level), our approach overcomes the challenge of excessively long input length for LLMs from long-form video. This enables LLMs to generate distinct types of questions, resulting in a fine-grained categorization aligned with the video content. To further enhance quality, a self-revision step refines results by removing redundancy and irrelevant prior information. For more details, please refer to section 2.
Duration Distribution.  The LongViTU dataset contains 1,833 videos, divided into 1,533 for training, 200 for validation, and 100 for testing, totaling approximately 900 hours. The average video duration is 29.3 minutes, ranging from 3.5 to 120.7 minutes with a standard deviation of 17.5 minutes, following a long-tail distribution (Figure 3(a)). Each video is accompanied by an average of 66 QA pairs, QA durations range from 6 to 1800 seconds, averaging 276.8 seconds, also exhibiting a long-tail pattern. Events and segments have average durations of 8.5 and 82 seconds, respectively. The dataset includes 121k QA pairs: 101k for training, 14k for validation, and 6k for testing.
Category Distribution.  The sunburst diagram in Figure 3(b) illustrates the distribution of question-answer pairs across three main categories: Spatiotemporal Understanding (55%), further divided into Object (12.2%), Attribute (10.7%), Location (15.5%), and Action (16.6%); Episodic Reasoning (24.4%), including Transition (8.1%), Interaction (3.4%), Causality (5.4%), and Motivation (7.5%); and Commonsense Inference (20.6%), comprising Planning (5.4%), Risk (2.7%), Function (6.4%), and Affordance (4.6%). Detailed categorization and examples are provided in Appendix E.
Computational Cost. We utilized 8 NVIDIA A100 GPUs (80GB) for dataset construction and all experiments. Dense captioning of all videos required approximately 1 week, while the generation pipeline consumed an average of 10k tokens per QA-pair via the OpenAI API [37]. Training session durations ranged from a minimum of 3 hours to a maximum of 3 days.
3 Experiments
We evaluated mainstream VLMs on the LongViTU test set, with instruction tuning conducted using our training set. The results demonstrate that LongViTU poses unique challenges for modern VLMs, whether through frame-based or sampling-based approaches. SFT significantly improved performance on both In-Distribution (ID) and several canonical Out-of-Distribution (OOD) benchmarks, highlighting the strong generalization and robustness of LongViTU.
(a) Video and QA duration distribution
(b) Sunburst of LongViTU
Figure 3: Statistics of LongViTU.  Subfigure (a) depicts the distributions of video and QA durations within LongViTU. The bottom horizontal axis (from left to right) represents QA duration in seconds, while the right vertical axis (from bottom to top) shows the percentage of the total dataset. QA durations predominantly vary from 6 to 600 seconds, with an average of 276.8 seconds. The top horizontal axis (from right to left) denotes video durations, and the left vertical axis (from top to bottom) indicates their percentage distribution. Video durations primarily range from 5 to 60 minutes, with an average of 29.3 minutes, following a long-tail distribution. Subfigure (b) illustrates the QA categorization in LongViTU along with their word frequency distribution. The outermost ring of the sunburst chart highlights the eight most frequent words in each category, with segment sizes reflecting their proportional frequencies. Please zoom in for a clearer view.
3.1 Setup
Settings and Baselines.  We fine-tuned multiple vision-language models (VLMs) on the complete LongViTU training set and evaluated their performance on the test set to address challenges in long-form video understanding, all input data was standardized in alignment with LLaVA [33]. Frame-based models such as mPLUG-OWL [61] and Video-LLaVA [31] uniformly sampled 8 frames per video as input. In contrast, LLaMA-VID [30], LongVA-DPO [65], LongVU [44], and Gemini-1.5-Pro [42] sampled 1 frame per second, while VideoAgent [15] sampled 1 frame every 2 seconds. All models, each with a uniform parameter size of 7B, generated answers to video-related questions. Evaluation employed a multi-level scoring system, where GPT-4 assessed the textual alignment between generated answers and ground truth based on predefined criteria, scores were averaged across subcategories to ensure a fair comparison.
Metrics and Benchmarks.  Traditional caption metrics are inadequate for evaluating open-ended questions. To address this limitation, we developed a multi-level scoring criteria that leverages GPT-4’s near-human capabilities in text comprehension and alignment. This approach rigorously evaluates the alignment between predicted answers and ground truth, ensuring that key aspects of each question are addressed. Hallucinations or irrelevant responses are penalized, while concise and accurate answers are rewarded with higher scores. Detailed scoring criteria and prompts are provided in Appendix C. Furthermore, all models were fine-tuned on the LongViTU training set and evaluated on the ID benchmark EgoSchema [36] and the OOD benchmarks VideoMME [16], WorldQA [67], and OpenEQA [35], achieving superior performance over baseline models.
3.2 Quantitative Analysis on LongViTU
The quantitative results on LongViTU detailed in subsection 3.2 reveal the following critical insights:
Effective Fine-Tuning.  Our supervised fine-tuning (SFT) led the open-source model LongVU to achieve a top GPT-4 score of $55.9$, surpassing the commercial Gemini-1.5-Pro’s $52.3$ at a dense sampling rate of 1 fps. This convergence between top open-source and commercial models highlights the challenges posed by LongViTU in facilitating comprehensive understanding of long-form video content.
Dense Sampling Challenge.  At a sampling rate of 1 fps, LLaMA-VID underperforms compared to mainstream open-source video language models in both zero-shot and fine-tuned settings. This gap reflects limitations in existing dense sampling strategies for capturing essential features in long-form video understanding.
LLM Reliance Bias.  The robust performance in text-only blind QA suggests a potential bias in QA systems favoring textual predictions. This phenomenon highlights over-reliance on textual intermediaries, allowing direct inference of answers from questions, further discussion in Appendix A.
[…]
Setting Method Overall Avg. Spatiotemporal Understanding
Object Attribute Location Action Avg.
Blind GPT-4 turbo 38.2 26.1 33.2 32.0 29.4 30.2
Frame-Based mPLUG-OWL 42.4 33.5 37.6 43.6 35.4 37.8
Video-LLaVA 45.9 37.8 46.3 49.1 38.1 42.7
Video-LLaVA∗∗ 50.7 39.3 49.2 49.6 41.8 44.9
[…]
Sampling-Based VideoAgent 44.0 35.7 43.1 45.9 36.4 40.2
LLaMA-VID 38.2 29.4 35.6 40.1 31.5 34.3
LLaMA-VID∗∗ 44.5 33.5 37.4 45.7 37.6 39.1
[…]
LongVA-DPO 47.5 35.9 52.4 44.3 37.2 41.8
LongVU 49.9 39.3 47.6 52.3 44.3 46.3
LongVU∗∗ 55.9 40.2 55.2 54.3 45.3 48.8
[…]
Gemini-1.5-Pro 52.3 54.3 58.6 56.3 48.1 54.7
Setting Method Overall Avg. Episodic Reasoning
Transition Interaction Causality Motivation Avg.
Blind GPT-4 turbo 38.2 45.1 47.4 47.7 56.1 49.5
Frame-Based mPLUG-OWL 42.4 45.8 47.7 47.7 49.4 47.6
Video-LLaVA 45.9 45.6 50.5 48.8 53.2 49.4
Video-LLaVA∗∗ 50.7 50.5 56.4 59.7 64.9 58.0
[…]
Sampling-Based VideoAgent 44.0 43.1 45.5 49.9 52.8 48.1
LLaMA-VID 38.2 40.4 46.7 40.5 46.6 43.2
LLaMA-VID∗∗ 44.5 46.7 48.4 54.2 57.7 52.1
[…]
LongVA-DPO 47.5 51.8 54.6 52.2 59.5 54.8
LongVU 49.9 56.7 58.8 48.9 53.4 54.2
LongVU∗∗ 55.9 62.4 63.5 63.2 70.0 65.2
[…]
Gemini-1.5-Pro 52.3 47.8 45.5 47.8 47.5 47.3
Setting Method Overall Avg. Commonsense Inference
Planning Risk Function Affordance Avg.
Blind GPT-4 turbo 38.2 36.5 51.1 55.9 50.9 48.7
Frame-Based mPLUG-OWL 42.4 42.1 54.6 54.3 51.5 50.3
Video-LLaVA 45.9 41.6 56.8 55.3 54.6 51.7
Video-LLaVA∗∗ 50.7 50.2 62.6 64.0 64.6 59.8
[…]
Sampling-Based VideoAgent 44.0 40.0 53.7 55.5 53.1 50.7
LLaMA-VID 38.2 34.9 51.3 46.5 47.2 44.1
LLaMA-VID∗∗ 44.5 43.9 54.5 55.7 53.8 51.7
[…]
LongVA-DPO 47.5 46.3 57.0 61.9 59.2 56.1
LongVU 49.9 46.2 52.3 60.4 62.8 56.0
LongVU∗∗ 55.9 59.9 67.9 68.7 70.4 66.4
[…]
Gemini-1.5-Pro 52.3 43.6 57.5 46.1 43.6 50.3
[…]
Method EgoSchema VideoMME WorldQA OpenEQA
Avg. Short Medium Long Avg. ScanNet HM3D
VideoLLM Online 47.4 13.7 24.3 16.7 0.0 30.0 23.3 24.8 20.4
[…]
LLaMA-VID∗∗ 34.0 17.2 23.8 15.4 12.2 32.2 33.6 33.5 33.8
[…]
Video-LLaVA 36.8 32.3 33.7 31.6 31.5 30.2 35.1 37.3 30.9
Video-LLaVA∗∗ 48.1 32.5 30.5 33.7 33.1 34.1 32.6 32.6 32.5
[…]
LongVA-DPO 56.9 54.3 61.6 50.4 47.6 30.3 36.6 41.5 26.9
LLaVA-OneVision 60.1 58.2 69.1 53.3 46.7 - - - -
LongVU 67.6 56.2 66.1 54.7 47.9 35.7 48.3 51.1 42.8
LongVU∗∗ 69.1 56.3 65.8 54.7 48.4 36.5 48.9 51.4 44.2
[…]
Spatial vs. Temporal Bias.  The weaker performance in Spatiotemporal Understanding reveals the unresolved challenges in long-form video comprehension. Tasks requiring detailed spatial analysis are particularly demanding. In contrast, Episodic Reasoning and Commonsense Inference leverage the logical sequence of events, resulting in improved performance through text-based methods.
Limitations of Commercial Models.  The commercial model Gemini-1.5-Pro demonstrates unexpectedly poor performance, marginally surpassing the zero-shot capabilities of open-source models while falling behind the fine-tuned open-source model LongVU. Due to resource constraints, full-length videos with specific timestamps embedded in the prompts were used instead of segmenting videos for each QA pair. Consequently, Gemini-1.5-Pro was required to process videos ranging from tens of minutes to over two hours for many questions, which significantly degraded its performance. These findings highlight the substantial challenges inherent in long-form video understanding.
3.3 Quantitative Analysis on Benchmarks
This section evaluates performance on benchmarks covering both In-Distribution (ID) and Out-of-Distribution (OOD) scenarios, leveraging datasets EgoSchema [36], VideoMME [16], WorldQA [67], and OpenEQA [35], as detailed in subsection 3.2. Key findings are outlined below.
Limitations of Sampling Models.  The LLaMA-VID [30] model exhibited a substantial performance drop during zero-shot evaluation in stage 3, which focuses on fine-tuning for long videos, compared to its performance in stage 2, which involves pre-training on images and short videos. This disparity underscores critical deficiencies in the stage 3 fine-tuning strategy. We revised the fine-tuning approach by applying all adjustments used in stage 2 as the baseline. The results demonstrated significant improvements across all benchmarks, with a maximum gain of 11.8% on EgoSchema.
Better on Longer Videos.  Fine-tuning with LongViTU resulted in decreased performance on shorter videos, as observed in the VideoMME Short subset and OpenEQA results. The average video duration of OpenEQA (49 seconds) is even shorter than that of the VideoMME Short subset (83 seconds). In contrast, performance on the Medium (563 seconds) and Long (2386 seconds) subsets improved. Specifically, LongViTU fine-tuning led to a 1% improvement on the VideoMME Long subset, while performance on the Short subset decreased and the Medium subset remained stable. This highlights the effectiveness of LongViTU in enhancing comprehension for longer videos, albeit with some trade-offs in performance on shorter ones.
Long-form Video Understanding Challenge.  The VideoLLM Online [7] model struggled significantly with the Long subset of VideoMME, producing incoherent responses and failing to generate measurable evaluations, underscoring the challenges of understanding long-form videos. The LongViTU dataset serves as both a vital pre-training resource and a benchmark for advancing long-form video understanding. This is demonstrated by the improved performance of models like LongVU on various additional benchmarks after supervised fine-tuning (SFT) on LongViTU, emphasizing the importance of optimizing models with large-scale high-quality video question-answering data.
3.4 Qualitative Evaluation
We present visualizations of various question-answering types in Figure 7 to facilitate a more thorough qualitative analysis. Due to space constraints, these figures and the ablation study have been deferred to Appendix B.
Spatial Details.  In Figure 7(a), the dense arrangement of multiple foreground objects in the scene led to incorrect zero-shot predictions by both Video-LLaVA and LongVU. Fine-tuning with LongViTU enabled the model to focus on finer spatial features, resulting in correct answers.
Key Moments.  As illustrated in Figure 7(b), the fine-tuned models accurately identified the key moment, a brief appearance of "a plant on the windowsill", and provided a precise and concise response, achieving a perfect GPT-4 score of 100. In contrast, the Video-LLaVA zero-shot model failed to capture this brief key moment.
Temporal Localization.  In Figure 7(c), both Video-LLaVA and LongVU correctly recognized "two" plug-in sockets in the kitchen at the end of a lengthy video and provided accurate responses after fine-tuning. However, both models failed in the zero-shot setting. This highlights the difficulty of extracting temporal details from extended video sequences and emphasizes the contribution of LongViTU data in improving generalization for long-form temporal localization.
3.5 Future Research
LongViTU introduces a novel direction in streaming video processing and video memory storage mechanisms for end-to-end models. Future research will build on these contributions, aiming to integrate streaming video understanding with question-answering, thereby paving the way for a more holistic approach to video understanding.
4 Related Work
Large Language Models.  Large language models (LLMs), such as InstructGPT [38], GPT-4 [1], LLaMA [48], and LLaMA-2 [49], have demonstrated notable capabilities in text processing and have been applied to the generation of large-scale datasets. These models convert different modalities into structured textual descriptions, which can then be used to prompt GPT-4 to produce multimodal data.
Instruction Tuning.  LLaVA [33] was among the first to leverage foundational vision models to generate image captions and detect bounding boxes, which were subsequently processed by ChatGPT or GPT-4 to generate image-based QA pairs. Building on this, methods like Bongard-OpenWorld [53], Video-LLaVA [31], and VideoChat [28] extended these principles to a more complex pipeline, shifting the focus from individual image QA to video-based QA. By sampling multiple frames from videos and applying LLaVA’s procedure, these approaches generate video QA datasets using structured frame descriptions, object categories, and attributes. However, the repetitive nature of frame-based descriptions, coupled with the input length limitations of LLMs, constrains the number of frames, thereby reducing the comprehensiveness of datasets.
Long-Context Language Models.  Even the most advanced long-context LLMs [46], such as GPT-4-turbo, ChatGLM [18], Baichuan2 [57], and InternLM2 [6], capable of handling input sequences beyond 128k tokens, suffer significant performance degradation when dealing with long and intricate texts. They struggle to manage the redundancy and disorganization in detailed video descriptions, limiting their effectiveness in generating video QA. Unlike static images, videos inherently demand an understanding of temporal dynamics, making event correlation crucial for video comprehension. The current frame-based approach fails to address this temporal aspect adequately, often resulting in shallow QA generation confined to individual frame analysis.
Long-Form Video Understanding.  Instruction tuning in the LLaVA paradigm [33] has demonstrated significant potential for multimodal tasks like captioning [8] and visual question-answering [4, 3, 47, 27, 12, 58, 2, 20, 39, 52]. Despite their effectiveness for images and short videos, these approaches encounter challenges with long-form videos [45, 31, 34, 63, 50, 65]. The primary limitation arises from the large number of visual tokens produced by encoders, which can range from 576 to 2880 tokens per image in models like LLaVA-NeXT [32]. As frame counts increase, these tokens rapidly surpass the context window limits of LLMs. To address this, recent methods use resamplers to reduce the visual tokens before passing them to LLMs [27, 30, 5, 10]. However, such reductions often degrade visual representations, resulting in diminished performance. Alternative approaches, including advanced pruning and feature merging, offer promising solutions [9, 43, 23, 68]. Notably, recent models such as VideoLLM Online [7], LongVILA [56], LLaVA-OV [26], LongVA [65], and LongVU [44] signify substantial advancements in long-form video understanding.
5 Conclusion
We present LongViTU, a large-scale dataset designed for long-form video understanding with integrated video memory and explicit timestamp annotations. Its hierarchical tree structure efficiently organizes video content, addressing critical challenges in generating QA pairs for extended videos. The self-revision process further enhances QA quality by mitigating hallucinations, redundancy, and irrelevant details. Fine-tuning on LongViTU demonstrates significant performance gains on both LongViTU and additional benchmarks, confirming its effectiveness and generalizability. Future work will explore advanced memory storage mechanisms that leverage explicit timestamps for advancing long-form video understanding and streaming question-answering.
Supplementary Material
Appendix A Limitation Statement
This paper focuses on the challenges of long-form video understanding and memory-based streaming video question-answering. Our automated pipeline demonstrates adaptability across diverse scenarios, with self-revision mechanisms that enhance data quality. However, some generated data may lack meaningful context, potentially impacting model performance. To mitigate this disadvantage, we propose refining the pipeline and incorporating manual inspection to better align the benchmark with real-world human annotations, and our work remains ongoing.
Appendix B Additional Results
Due to space limitations, we defer more details about the ablation study, human study, and qualitative evaluation to this section.
B.1 Ablation Study
Varying Durations.  Figure Figure 4 illustrates LongVU’s performance on LongViTU with varying QA durations. As the QA duration (in seconds) increases, the performance gradually declines, consistent with the trends observed in VideoMME across all baseline models in subsection 3.2.
Figure 4: Duration Ablation.  After fine-tuned on LongViTU, LongVU’s performance exhibits a mitigated downward trend on longer subsets [300, 600) compared to shorter subsets.
Training Epochs.  Figure Figure 5 illustrates LongVU’s performance on Egoschema improves after SFT as the training epochs derived from LongViTU increase. The performance gains diminish as the training epochs approach the full scale of LongViTU, eventually plateauing at its optimal performance.
Figure 5: Epoch Ablation.  Performance improvements during fine-tuning diminish as training epochs approach the full scale of LongViTU.
B.2 Human Study
We randomly selected 100 QA pairs from both LongViTU and VideoMME for a human study conducted by three evaluators, as shown in Figure 6.
Self-Evaluation.  Each QA pair was assessed for its alignment with the video content. Good pairs were concise, accurate, and spatiotemporally relevant with video content. Fair pairs matched the video but lacked spatiotemporal relevance. Poor pairs included irrelevant or inaccurate information.
LongViTU vs. VideoMME.  QA pairs were assessed for coherence, and alignment with video content, comparing their similarity to human-crafted pairs. The results were categorized as: VideoMME significantly better, LongViTU significantly better, or Tie for comparable quality.
Figure 6: Human Preference.  In self-evaluation, approximately 91% of QA pairs were marked as correct. In comparison, around 64% of QA pairs were comparable in quality to VideoMME.
B.3 Qualitative Evaluation
A detailed qualitative analysis is provided in subsection 3.4 and illustrated in Figure 7.
(a) Spatial Details.  The dense arrangement of multiple foreground objects in the scene led to incorrect zero-shot predictions by both Video-LLaVA and LongVU. Fine-tuning with LongViTU enabled the model to focus on finer spatial features, resulting in correct answers.
(b) Key Moments.  The fine-tuned models accurately identified the key moment, a brief appearance of "a plant on the windowsill", and provided a precise and concise response, achieving a perfect GPT-4 score of 100. In contrast, the Video-LLaVA zero-shot model failed to capture this brief key moment.
(c) Temporal Localization.  Both Video-LLaVA and LongVU correctly recognized "two" plug-in sockets in the kitchen at the end of a lengthy video and provided accurate responses after fine-tuning. However, both models failed in the zero-shot setting. This highlights the difficulty of extracting temporal details from extended video sequences and emphasizes the contribution of LongViTU data in improving generalization for long-form temporal localization.
Figure 7: Qualitative Results.  The yellow box indicates the key frame that contains the answer, while the red box highlights the relevant objects. For better illustration, only concise key information is presented in the predictions.
Appendix C More Details on Building LongViTU
C.1 Hierarchical Video Tree Construction
This subsection outlines the hierarchical video tree construction process, with the details provided in subsubsection 2.1.1, and the algorithm elaborated in Algorithm 1. The corresponding prompts for each stage are described in the following.
Frame Level.  For dense captioning at the frame level, the "internlm-xcomposer2-vl-7b-4bit" model is employed with the following prompt:
<ImageHere>Identify each object in the image, describe their positions, and detail their appearance.
Event Level.  We employ the ChatCompletion API of the "gpt-4-turbo-2024-04-09" model with the following prompt to refine event level descriptions:
Write a concise narrative in one sentence, including visual details from "Frames" that depict an "Event", do not use any unrelated information.
"Event" describes an action in a video, with "C" representing me and other letters like ’X’ and ’Y’ standing for different people, transform these for a smoother narrative.
"Frames" show detailed visuals and space details of objects in each moment during the "Event".
Event: {event}
Frames: {frames}
Just return narrative that summarizes the episodic memory depicted in this video, only focuses on spatial details and temporal correlations.
Narrative:
Segment Level.  We utilize the ChatCompletion API of the "gpt-4-turbo-2024-04-09" model to generate segment level descriptions:
Integrate sequential event descriptions of video content into a very concise summary in one sentence, from my perspective for a smoother narrative. Each segment should capture a sequence of closely related actions, events, or scenes. Using "index" to represent the start and end of each segment, do not use any unrelated information.
0:   Annotated events $Events$ in Ego4D
0:   Hierarchical video tree $\mathcal{T}_{\text{video}}$
11:   Initialize $\mathcal{T}_{\text{video}}=\emptyset$;
22:   Sample frames $Frames$ at 1 fps across annotated events;
3:   for each frame $f$ in $Frames$ do
34:      $d_{f}=\text{InternLM-XComposer2}(f)$;
45:      $t_{s}^{f}=f.\text{timestamp}$; $t_{e}^{f}=f.\text{timestamp}$;
56:      $Frames[f]=\langle d_{f},t_{s}^{f},t_{e}^{f}\rangle$;
7:   end for
8:   for each event $e$ in $Events$ do
69:      Collect frames $\{\langle d_{f},t_{s}^{f},t_{e}^{f}\rangle\}$ within event $e$;
710:      $d_{e}=\text{GPT-4}(\{d_{f}\})$;
811:      $t_{s}^{e}=e.\text{start}$; $t_{e}^{e}=e.\text{end}$;
912:      $Events[e]=\langle d_{e},t_{s}^{e},t_{e}^{e},\{\langle d_{f},t_{s}^{f},t_{e}^{f}\rangle\}\rangle$;
13:   end for
1014:   Group related events into segments $\{Segments\}$ using GPT-4;
15:   for each segment $s$ in $Segments$ do
1116:      $d_{s}=\text{GPT-4}(\{d_{e}\text{ in }s\})$;
1217:      $t_{s}^{s}=\min\{t_{s}^{e}\text{ in }s\}$; $t_{e}^{s}=\max\{t_{e}^{e}\text{ in }s\}$;
1318:      $\mathcal{S}[s]=\langle d_{s},t_{s}^{s},t_{e}^{s},\{Events\text{ in }s\}\rangle$;
19:   end for
1420:   $\mathcal{T}_{\text{video}}=\langle R,\{\mathcal{S}[s]\}\rangle$;
1521:   return  $\mathcal{T}_{\text{video}}$;
Algorithm 1 Hierarchical Video Tree Construction
Step-by-step:
1. Review event descriptions and group consecutive events that are closely related into a segment.
2. For each group of events, write a brief summary.
"index" represents order of event, "event" outlines this moment.
Video Content:
{video content}
Return each segment in JSON format: "start": start index, "end": end index, "segment": brief description of video segment. Assemble all segments into a single Python list, ensuring output is neatly organized and strictly adheres to this JSON format.
Segments:
C.2 Long-Form QA Generation
QA Generation.  We adopt the ChatCompletion API of the "gpt-4-turbo-2024-04-09" model to generate QA pairs on the selected subtree:
Task:
Construct episodic memory of video content through question-answer pairs that encapsulate spatial and temporal aspects within selected events.
Step-by-Step Instructions:
1. Selection of Events: Select either a single specific event or a series of interrelated events from the video content (’Memory Content’). For each selected event or sequence of events, generate question-answer pairs that reflect their spatial and temporal characteristics. Use "index" to designate the chronological order of these memory events. 2. Creation of Question-Answer Pairs: From the selected events, formulate questions that will be posed later in the video related to a single, specific event (’Ask Content’). These pairs should mimic a retrospective dialogue between me and an AI assistant, where I pose questions and the AI provides answers based on the video content. Reference events and segments to make dialogue more naturally narrative, avoiding direct references "index" or timestamps.
3. Categorization of Questions: Categorize each question under a specific type such as: Object, Attribute, Location, Action, Function, Affordance, Comparison, Relationship, Causality, Motivation, Planning, Risk, or any other category you suggest.
Output Format:
Return question-answer pairs in JSON format: "memory": [list of memory events index], "ask": event index where question is posed, "type": question type, "question": question, "answer": answer. Assemble all pairs into a single Python list, ensuring the output is neatly organized and strictly adheres to this JSON format.
Term Definitions of Video Content:
- segment: a brief summary covering a sequence of related events.
- events: multiple related events within a segment.
- index: sequential position of an event within the overall video content.
- event: spatial-temporal details associated with each moment in the video.
Memory Content:
{memory content}
Ask Content:
{ask content}
Question-Answer pairs:
C.3 Self-Revision
Self-Revision.  We utilize the ChatCompletion API of the "gpt-4-turbo-2024-04-09" model to perform self-revision:
Please review and correct the following question-answer pair about video content. Simplify the pair to directly represent the core information without redundant details, ensuring the question is natural and concise, and the answer is direct and clear. Identify the correct type of the QA pair: Object, Attribute, Location, Action, Function, Affordance, Comparison, Relationship, Causality, Motivation, Planning, Risk, or Other. Do not add or fabricate content. Remove redundant event numbers and express the event directly.
Original QA:
{original qa}
Return the Revised QA as a dict:
{"revised type": revised QA type, "revised question": revised question, "revised answer": revised answer}
Revised QA:
C.4 Evaluation Metrics
Scoring Criteria.  We use the ChatCompletion API of the "gpt-4-turbo-2024-04-09" model to perform evaluation by designed scoring criteria:
As a scoring expert, your responsibility is to evaluate the accuracy of a model’s response to a specific question about video content. You will be provided with the ’question’ asked about the video, the ’answer’ which is the correct answer based on the video, and the ’prediction’ which is the model’s response. Your task is to assess how accurately the model’s ’prediction’ answers the ’question’ in relation to the ’answer’.
Question:
{question}
Answer:
{answer}
Prediction:
{prediction}
Scoring Criteria:
Level 1: The ’prediction’ is unrelated to the ’question’ or unintelligible, containing significant errors or irrelevant characters. Score: 0.
Level 2: The ’prediction’ is completely off-topic, not reflecting the factual content of the ’answer’. Score: 20.
Level 3: The ’prediction’ somewhat response the ’question’ but includes errors or irrelevant details not found in the ’answer’. Score: 40. Level 4: The ’prediction’ generally response the ’question’ but has some inaccuracies or irrelevant details compared to the ’answer’. Score: 60.
Level 5: The ’prediction’ accurately response the ’question’ and is mostly consistent with the ’answer’, with only minor discrepancies. Score: 80.
Level 6: The ’prediction’ perfectly response the ’question’ and fully aligns with the facts provided in the ’answer’. Score: 100.
Only provide the numerical score based on the criteria above without any additional commentary.
Score:
Appendix D Implementations and Hyperparameters
LABEL:tab:params presents the key training parameters of our primary fine-tuning models: LLaMA-VID, Video LLAVA, and LongVU. Pretraining and implementation details for these models are described below.
LLaMA-VID integrates the EVA-G visual encoder with the QFormer text decoder. Key strategies include freezing both the visual encoder and the text decoder during all stages, except for the uninitialized BERT module, which is optimized for one epoch per stage.
Video-LLaVA processes input images by resizing and cropping them to 224×224, with 8 frames sampled per video. Training is conducted in two stages: the first stage employs a batch size of 256, while the second stage reduces it to 128.
LongVU utilizes SigLIP and DINOv2 for visual encoding, combined with Qwen2-7B for language processing. Training involves cross-entropy loss for text generation, optimized using the AdamW optimizer with a cosine schedule. Image-language pretraining is performed with a batch size of 128 and a learning rate of 1e-5. During video-language fine-tuning, the batch size is reduced to 64, with adaptive compression techniques applied. The model employs sliding windows and spatio-temporal compression methods for efficient processing.
Table 4: Fine-tuning Hyperparameters.  This table details the primary training parameters utilized for fine-tuning LLaMA-VID, Video-LLaVA, and LongVU. Parameters not explicitly listed here adhere to the default configurations of their respective implementations.
Parameter Value
LLaMA-VID
$Pretrained~{}LLM$ $Vicuna\_1.5\_7B$
$max\_length$ $2048$
$batch\_size$ $1$
$gradient\_accumulation$ $64$
$learning\_rate$ $5e^{-6}$
$weight\_decay$ 00
$warmup\_ratio$ $0.03$
$lr\_scheduler$ $cosine$
$image\_token$ $2$
$video\_fps$ $1$
Video-LLaVA
$Pretrained~{}LLM$ $LLaMA\_2\_7B$
$max\_length$ $2048$
$batch\_size$ $32$
$gradient\_accumulation$ $2$
$learning\_rate$ $1e^{-5}$
$weight\_decay$ 00
$warmup\_ratio$ $0.03$
$lr\_scheduler$ $cosine$
$image\_token$ $256$
$video\_frames$ $8$
LongVU
$Pretrained~{}LLM$ $Qwen\_2\_7B$
$max\_length$ $8192$
$batch\_size$ $32$
$gradient\_accumulation$ $1$
$learning\_rate$ $5e^{-7}$
$weight\_decay$ 00
$warmup\_ratio$ $0.03$
$lr\_scheduler$ $cosine$
$image\_token$ $144$
$video\_fps$ $1$
Appendix E More LongViTU Examples
Table 5: Examples of LongViTU.  We demonstrate the ratio of each concept category and more examples of LongViTU. For a more intuitive perspective, you may refer to Figure 3.
Category Ratio Question Example (Q) Answer Example (A)
Object 12.2% What am I holding in my hand? A cup.
What items are on the table? A pen and a notebook.
Attribute 10.7% What is the color of that clothing? Blue.
What is the material of the cup? Porcelain.
Location 15.5% Where am I in the house right now? In the kitchen.
Where is the key placed in? Inside the drawer.
Action 16.6% What is that man doing? He is cooking.
What am I doing by the counter? Preparing coffee.
Transition 8.1% Where did he go after leaving here? To the garden.
What change happened to that cup? It fell and broke.
Interaction 3.4% Which hand did I use to pick up this wrench? The left hand.
What did I take out of the microwave? A bowl of soup.
Causality 5.4% What happened after I pressed that button? The light turned on.
What happened after I opened the box? A toy popped out.
Motivation 7.5% Why should I leave the room? There is smoke.
Why does she want to open the cabinet? To get a plate.
Planning 5.4% How do I get to the backyard? Through the kitchen door.
How do I repair this house? By fixing the roof first.
Risk 2.7% What dangers does that saw pose? It could cut your hand.
What dangers are there in the kitchen? Slippery floor.
Function 6.4% What is the function of this tool? To tighten bolts.
What is the function of this box? To store items.
Affordance 4.6% What can this stone be used for? To grind spices.
What can this glass bottle be used for? To store water.